OpenStudy (anonymous):

If z being a complex number lies on the unit circle and the sum of z with its reciprocal is 1, then find z^235

OpenStudy (anonymous):

wt bt this??

OpenStudy (amistre64):

z^235? thats alot

OpenStudy (anonymous):

theres no answer given bt frm what i know it comes out to be a simple answer

OpenStudy (amistre64):

z + 1/z = 1?

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

z^2 +1 ------- = 1; z^2 +1 = z; z^2 -2 +1 = 0 Z

OpenStudy (amistre64):

(z-1)^2; z = 1

OpenStudy (amistre64):

z^235 = 1 then ?

OpenStudy (anonymous):

but z isnt 1, its a complex number...

OpenStudy (amistre64):

typoed it again

OpenStudy (anonymous):

z is the 3rd root of -1. because z^2-Z+1=0; z^3=1 z^235=z=-1+isqrt(3) /2

OpenStudy (amistre64):

z^2+1 = 1 z^2 = 0 z=0 ?

OpenStudy (anonymous):

Z^3=-1 (z+1)(z^2-z+1)=0 Z^2-z+1=0;

OpenStudy (amistre64):

so that means the the complex part is either up or down

OpenStudy (anonymous):

z+1/z=1 z^2-z+1=0.

OpenStudy (anonymous):

i think dipankars right, did u apply the demoivres theorem??

OpenStudy (anonymous):

no need to apply it....

OpenStudy (amistre64):

dip is usually right ;)

OpenStudy (anonymous):

i have given the steps....

OpenStudy (anonymous):

z is the complex roots of -1.

OpenStudy (anonymous):

so but then howdu get z^235??

OpenStudy (anonymous):

u must realize amistre m doing wht u often do;)

OpenStudy (amistre64):

guess wildly lol

OpenStudy (anonymous):

z^235=z^234*z=(z^3*78)^z=1*z=z

OpenStudy (anonymous):

no not that...like wht u were doin with the tangent plane question...

OpenStudy (anonymous):

ok dipankar....

OpenStudy (amistre64):

ahhh... :)

OpenStudy (anonymous):

with all the three ;)

OpenStudy (amistre64):

its good to be smart :)

OpenStudy (anonymous):

him i think u have understood the process...

OpenStudy (anonymous):

yeah right...

OpenStudy (amistre64):

so, when I find z=0, the complex parts have to be pure imaginaries then right? like "i"

OpenStudy (anonymous):

ill tell u what

OpenStudy (anonymous):

z + 1/z =1 let z=cosθ + isinθ 1/z = cosθ - isinθ z + 1/z = 2cosθ = 1 cosθ = 1/2 hence z = 1/2 + √3/2

OpenStudy (amistre64):

z^235 = i^235...235/4 = 58 and 3 remaining i, -1, -1 z^235 = -i?

OpenStudy (anonymous):

√3/2 x i sorry

OpenStudy (anonymous):

no man look above z = 1/2 + i√3/2

OpenStudy (amistre64):

lol...stab in the dark then :) i see what you did

OpenStudy (anonymous):

now du wanna know the other one as well?

OpenStudy (amistre64):

i do, but at this late hour it aint gonna stick to well ;)

OpenStudy (anonymous):

hahaha...wts d time there??